In the game of roulette, a metal ball is spun around a rotating wheel containing 18 red-numbered slots, 18 black-numbered slots, and 2 green slots numbered 0 and 00, respectively. Find the probability that the ball falls into a red slot exactly 14 times in 20 spins. 0.7000 0.0370 0.4737 0.0236
Answer #1

P( red) = p=18/38=0.4736

P( red=14)= 20C14 * 0.4736^14 * (0.5264)^6=20!/[14!*6!]* 0.4736^14 * (0.5264)^6=0.02355 or 0.0236